# AirRadar: Inferring Nationwide Air Quality in China with Deep Neural Networks

### v1

@@ Introduction
Targeting the first challenge, we propose a Spatial Learning Module that captures spatial correlations locally and globally. The local learner utilizes dartboard projection (Liang et al. 2023) to effectively process the local information. The global learner uses the Fast Fourier Transform (FFT) to transform the feature into the frequency domain and capture global relationships with less computational complexity. For the second challenge, our Casual Learning Module applies causal theory to interpret causal-effect relationships in real-world contexts, dynamically weighting each context based on its contribution.
@@
Casual Inference. Causal inference (Pearl, Glymour, and Jewell 2016) aims to explore causal relationships between variables, promoting stable and robust learning and inference processes. The integration of deep learning techniques with causal inference has demonstrated significant potential in spatial-temporal mining tasks (Xia et al. 2024; Luo et al. 2024). In this study, we apply causal methods to address spatial heterogeneity in the air quality inference task.
@@
  • Casual Learning Module: We begin by constructing a structural causal model to illustrate how various contexts influence air quality. We then implement backdoor adjustment by dynamically allocating weights to each context, effectively mitigating confounding effects.
@@
Casual Learning Module
@@
Estimating this distribution is facilitated by the Casual Learning Module, illustrated in Figure 3c. Assuming that the contexts are independent can lead to overfitting issues, thus we denote a context as a 0-1 matrix, i.e.,
@@
Our model, implemented with PyTorch 2.1.0 on an NVIDIA RTX 2080Ti GPU, uses the Adam optimizer with a batch size 32, and an initial learning rate of $5\times 10^{-3}$. The hidden size ($E$) is 32. We randomly mask a percentage of node features and include 2 spatial learning blocks. The Local part has 2 heads with 50km and 200km ranges. The global part comprises 2 weight blocks, and the Casual Learning Module uses 4 contexts.
@@
Effects of Casual Learning Module. To evaluate the effectiveness of the Casual Learning Module in addressing spatial heterogeneity challenges, we compare our model with the e following variants: a) w/o Causal: The entire Casual Learning Module is omitted from our AirRadar. b) w/o Adaptive: The dynamic weight allocation mechanism is excluded from the Casual Learning Module. The outcomes are presented in Figure 7. When the causal module is omitted, the model fails to mitigate the influence of potential context and, consequently, cannot address the spatial heterogeneity issue. Absent the adaptive module, the model is unable to account for varying impacts introduced by different contexts, resulting in a performance decline.

### v2

@@ Introduction
Targeting the first challenge, we propose a Spatial Learning Module that captures spatial correlations locally and globally. The local learner utilizes dartboard projection (Liang et al. 2023) to effectively process the local information. The global learner uses the Fast Fourier Transform (FFT) to transform the feature into the frequency domain and capture global relationships with less computational complexity. For the second challenge, our Causal Learning Module applies causal theory to interpret causal-effect relationships in real-world contexts, dynamically weighting each context based on its contribution.
@@
Causal Inference. Causal inference (Pearl, Glymour, and Jewell 2016) aims to explore causal relationships between variables, promoting stable and robust learning and inference processes. The integration of deep learning techniques with causal inference has demonstrated significant potential in spatial-temporal mining tasks (Xia et al. 2024; Luo et al. 2024). In this study, we apply causal methods to address spatial heterogeneity in the air quality inference task.
@@
  • Causal Learning Module: We begin by constructing a structural causal model to illustrate how various contexts influence air quality. We then implement backdoor adjustment by dynamically allocating weights to each context, effectively mitigating confounding effects.
@@
Causal Learning Module
@@
Estimating this distribution is facilitated by the Causal Learning Module, illustrated in Figure 3c. Assuming that the contexts are independent can lead to overfitting issues, thus we denote a context as a 0-1 matrix, i.e.,
@@
Our model, implemented with PyTorch 2.1.0 on an NVIDIA RTX 2080Ti GPU, uses the Adam optimizer with a batch size 32, and an initial learning rate of $5\times 10^{-3}$. The hidden size ($E$) is 32. We randomly mask a percentage of node features and include 2 spatial learning blocks. The Local part has 2 heads with 50km and 200km ranges. The global part comprises 2 weight blocks, and the Causal Learning Module uses 4 contexts.
@@
Effects of Causal Learning Module. To evaluate the effectiveness of the Causal Learning Module in addressing spatial heterogeneity challenges, we compare our model with the e following variants: a) w/o Causal: The entire Causal Learning Module is omitted from our AirRadar. b) w/o Adaptive: The dynamic weight allocation mechanism is excluded from the Causal Learning Module. The outcomes are presented in Figure 7. When the causal module is omitted, the model fails to mitigate the influence of potential context and, consequently, cannot address the spatial heterogeneity issue. Absent the adaptive module, the model is unable to account for varying impacts introduced by different contexts, resulting in a performance decline.